# Middo: Model-Informed Dynamic Data Optimization for Enhanced LLM Fine-Tuning via Closed-Loop Learning

### v1

@@ 5 Analysis
iteration Ablations IFEval MATH HumanEval Hellaswag Average
@@
5.1 Ablation Studies
To assess the effectiveness of Middo and the contribution of each optimization pipeline, we conduct ablation experiments on the LLaMA-3.1-8B model over the Alpaca dataset. Specifically, we analyze the following ablations: (a) w/o loss: removes Loss Patterns. (b) w/o neighbor: excludes Embedding Cluster Dynamics. (c) w/o score: removes Self-alignment Scores.
@@ 5.2 Effect of Selected Data Scale
We give an investigation on the impact of the different scales of the selected and optimized data in this section by varying the thresholds for data selection. Results are illustrated in Figure 3. We observe that increasing the size of the refined data initially leads to an upward trend in performance; however, once the refined data exceeds a certain threshold, performance begins to decline. To maintain the potential for further iterative improvement, we set the refined data size at a moderate level that optimally balances the cost and benefit of the optimization process. In the first iteration, each component selects approximately $5\%$ of the data for refinement. By controlling the parameter $m$, the amount of data refined can adaptively change as the model’s capability increases. Detailed data sizes selected in each iteration are provided in Appendix E.
@@ 5.3 Data Analysis
For a deep understanding of how Middo transforms the dataset, we provide an analysis of its impact on data complexity, diversity, and quality.
@@
Figure 4: Loss distribution comparison before and after applying Middo. We tested LLaMA-3.1-8B on the Alpaca dataset over one optimization iteration. The density curve reflects the relative frequency of data points within specific loss intervals. The inset subfigure highlights the maximum loss reduction from 12.99 to 4.61 through our optimization.
To quantify how Middo modulates dataset complexity, we analyze the loss distribution evolution through optimization cycles. As shown in Figure 4, the original dataset exhibits a long-tailed distribution with extreme loss values up to $12.99$. After applying Middo, the maximum loss decreases by $64.5\%$ to $4.61$, indicating successful mitigation of overly complex samples and the distribution mode shifts leftward, suggesting better alignment between data complexity and model capability. This transformation demonstrates our framework’s ability to adaptively prune pathological samples while preserving pedagogically valuable challenges.
@@
Figure 5: t-SNE visualization of the Alpaca dataset before and after applying Middo. The original dataset is shown in light blue, while the augmented data is in dark blue. The plot illustrates the distribution of data points in the latent space, highlighting the impact of Middo on dataset diversity.
To analyze the diversity of the dataset after applying Middo, we visualize the data distribution using t-SNE (Van der Maaten and Hinton, 2008). Figure 5 reveals how the augmented data points are distributed relative to the original data. Notably, most of the augmented samples are located at the peripheries of the clusters, effectively filling in the sparsely populated regions. This distribution indicates that Middo is not merely adding redundant data but is instead enhancing the overall coverage of the latent space. By strategically augmenting the dataset at the cluster edges, Middo improves the diversity and ensures a more uniform distribution of data points, ultimately contributing to better model generalization.
@@ 6 Conclusion
In this paper, we present Middo, a model-informed dynamic data optimization framework that transforms LLM fine-tuning via closed-loop learning. Unlike traditional static methods, Middo establishes a self-evolving system that continuously adapts to the model’s evolving capabilities. It employs three core mechanisms: complexity optimization refines overly complex samples using loss patterns, ensuring the training data remains appropriately challenging; diversity optimization enhances dataset diversity by analyzing embedding cluster dynamics and addressing gaps in the semantic space; and quality optimization leverages self-alignment scores to evaluate and improve the quality of training samples. Experiments on multiple benchmarks demonstrate that Middo consistently boosts LLMs’ performance, achieving an average accuracy improvement of $7.15\%$ while maintaining the original data scale on LLaMA-3.1-8B. Ablation studies confirm the effectiveness of each component, underscoring the importance of balancing complexity, diversity, and quality. Middo ’s adaptability and model-awareness make it a powerful tool for sustainable LLM training, ushering in a new paradigm of dynamic human-AI co-evolution of data and models. Moreover, our approach paves the way for future research in adaptive training that continuously optimize learning efficiency.
@@
Appendix A Computational Cost Analysis
We analyzed the computational cost of Middo’s optimization stages on 7B parameter models (LLaMA-3.1-8B, Mistral-7B-v0.3) using 50k-100k sample datasets (Alpaca, Alpaca-4o-mini, WizardLM) on 8 $\times$ NVIDIA A100 GPUs.
Each optimization iteration, encompassing data selection via loss patterns, embedding cluster dynamics, and self-alignment scores, followed by refinement, typically completes in under 30 minutes. This efficiency is largely due to the parallelizable nature of the diagnostic modules and the use of acceleration techniques: CUDA for neighbor computation in Embedding Cluster Dynamics and vLLM (Kwon et al., 2023) for batched inference during Self-alignment Score calculation. Table 4 provides a detailed time breakdown per component, underscoring Middo’s practical efficiency.
Method Component Time (Single A100 GPU) Time (8 $\times$ A100 GPUs, Data Parallelism)
Loss patterns < 50 minutes < 10 minutes
Embedding cluster dynamics < 40 minutes + neighbor computation time < 10 minutes (CUDA acceleration)
Self-alignment scores < 1 hour per metric (6 metrics) < 10 minutes (vLLM acceleration)
Table 4: Approximate overhead time cost of Middo’s optimization components per iteration. Timings are reported for processing datasets in the range of 50k-100k samples.
Appendix B Hyperparameters Analysis
B.1 The Impact of Neighbor Number
$\bm{k}$ IFEval GSM8K MATH HumanEval MBPP Hellaswag ARC-c Average
11 43.59 38.74 9.20 35.98 39.8 48.59 17.17 33.3
22 51.56 43.21 10.72 40.85 41.00 57.47 12.12 35.72
33 40.82 40.49 9.50 32.32 39.20 59.72 8.59 32.95
Table 5: Impact of the number of neighbors ($k$) in the Embedding Cluster Dynamics on Middo performance. The table shows the performance across various benchmarks for different values of $k$, indicating that $k=2$ yields the best overall average score.
We also explore how the number of neighbors $k$ used in the Embedding Cluster Dynamics affects the overall performance of Middo. By varying the number of neighbors, we analyze its impact on dataset diversity and model performance. Table 5presents the results of this analysis. We find that the optimal number of neighbors is $k=2$, which achieves the best balance between diversity and performance. This setting ensures that the dataset is sufficiently expanded to enhance model generalization while avoiding excessive noise that may degrade performance.
B.2 The Impact of Iterations
We tested the number of iterations on the Alpaca dataset as Figure 7 shows and found that the model’s performance significantly declined after the third iteration. Therefore, we chose to optimize each dataset for three iterations. This optimal number is not necessarily fixed and may vary depending on the threshold of each iteration.
B.3 The Impact of Thresholds
Dataset iteration loss neighbor self total
LLaMA-3.1-8B
Alpaca init $m=1$ m=1m=-1 m=1.5m=-1.5 52,002
iter1 1180 1924 1159 53,939
iter2 299 1853 108 55,811
iter3 242 1822 381 57,636
Alpaca 4o-mini init $m=0$ m=1m=-1 m=0.5m=-0.5 52,002
iter1 5684 8032 4145 60,865
iter2 611 2291 876 63,184
iter3 472 2127 661 65,324
Wizard init $m=1$ m=1.5m=-1.5 m=2m=-2 70,000
iter1 3585 3585 2690 73,642
iter2 959 3341 1016 76,993
iter3 751 3414 420 80,419
Mistral-7B-v0.3
Alpaca init $m=0.5$ m=2m=-2 m=1m=-1 52,002
iter1 2418 2111 2367 54,131
iter2 1985 2091 932 56,268
iter3 1788 1982 352 58,348
Alpaca 4o-mini init $m=1$ m=2m=-2 m=2.5m=-2.5 52,002
iter1 1407 7691 1499 59,696
iter2 1278 9116 1045 68,874
iter3 1346 2487 661 74,036
Wizard init $m=1$ m=1.5m=-1.5 m=1.5m=-1.5 70,000
iter1 5637 5709 5258 76,429
iter2 3558 5999 6310 82,501
iter3 3885 6229 3767 89,178
Table 6: Data Size Details Across Iterative Refinement. For each dataset, the table lists the number of samples selected by the three components—loss (Loss Patterns), neighbor (Embedding Cluster Dynamics), and self (Self-alignment Scores). During each iteration, along with the total data size after refinement. The init row represents the original dataset size and the threshold controlling hyperparameter $m$ corresponding to each component.
The amount of data selected for refinement by each module (Loss Patterns (Complexity), Embedding Cluster Dynamics (Diversity), and Self-alignment Scores (Quality)) is governed by dynamic thresholds $\tau=\mu+m\sigma$, where $\mu$ and $\sigma$ are the mean and standard deviation of the respective signal values (loss, diversity score, quality score) across the dataset. The multiplier $m$ is a key hyperparameter that controls the stringency of these thresholds.
Figure 7: Performance trends on the Alpaca dataset across different iterations. The model’s performance peaks at three iterations and declines thereafter.
Our approach to setting $m$ is guided by empirical analysis aimed at optimizing refinement effectiveness. Initial experiments (detailed in Section 5.3, Figure 3) indicated that refining a total unique proportion of approximately $10-20\%$ of the dataset in the first iteration yields substantial performance improvements.
To determine appropriate $m$ values for each module, we conducted a sensitivity analysis, presented in Table 7. This table shows how different combinations of $m$ for complexity, diversity, and quality impact the total percentage of unique data selected for refinement and the resulting average model performance on benchmarks. The $m$ values are varied (e.g., in $0.5$ increments) for each signal, and combinations are chosen to target the 10-20% total selected data range. As shown, performance peaks when the selected data proportion falls within this empirically determined optimal range. For instance, the combination yielding $14.88\%$ selected data achieved the best average score of $43.23$. When multiple $m$ combinations meet the $10-20\%$ criterion, we select those with the smallest absolute $m$ values (representing the mildest effective thresholds) that achieve this target, balancing refinement impact with efficiency.
Complexity Diversity Quality Total Selected Percentage Performance
$m=0$ m=1m=-1 m=1.5m=-1.5 15.8k 30.45% 41.81
$m=0.5$ m=1.5m=-1.5 m=1.5m=-1.5 7.7k 14.88% 43.23
$m=1$ m=2m=-2 m=1.5m=-1.5 4.3k 8.20% 41.96
$m=1.5$ m=2.5m=-2.5 m=1.5m=-1.5 2.6k 5.09% 41.55
$m=2$ m=3m=-3 m=4m=-4 1.3k 2.44% 40.87
$m=3$ m=3.5m=-3.5 m=10m=-10 0.5k 0.92% 39.64
$m=4$ m=4m=-4 m=12m=-12 0.1k 0.26% 38.69
Table 7: Sensitivity analysis for the threshold multiplier $m$ on the Alpaca dataset (first iteration). The table shows the impact of varying $m$ for complexity, diversity, and quality modules on the total unique data selected (sum and percentage) and the average model performance (mean score across benchmarks).
The actual data sizes selected in each iteration for the experiments reported in the main paper, using $m$ values derived from this sensitivity analysis (e.g., targeting the $15\%$ mark initially), are detailed in Table 6. As the model’s performance improves over subsequent iterations, the amount of data flagged by these fixed $m$ thresholds naturally decreases due to shifts in the signal distributions ($\mu$ and $\sigma$). This adaptive selection aligns with our observation that early training phases benefit from addressing a broader set of initial complexities and diversities, while later stages refine more nuanced aspects.
We do not place excessive emphasis on the improvements brought about by differences in data volume, so our selection may not necessarily be optimal.
Appendix C Experimental Details
C.1 Instruction Fine-tune Dataset
We evaluate Middo on three general instruction fine-tuning datasets.
  • Alpaca (Taori et al., 2023): consists of 52,002 instruction-response pairs generated by Stanford University using the self-instruct (Wang et al., 2023) method based on OpenAI’s text-davinci-003. This dataset is designed for fine-tuning dialogue models similar to ChatGPT to achieve efficient instruction-following capabilities.
  • Alpaca-4o-mini: to evaluate performance on a higher-quality response dataset, we generated responses for all Alpaca instructions using GPT-4o mini, creating the Alpaca-4o-mini dataset.
  • WizardLM (Xu et al., 2024): 70K data generated based on Evol-Instruct, which aims to generate more complex instruction data through a recursive evolutionary approach in order to improve the model’s reasoning and instruction comprehension.
C.2 Models
We primarily conducted experiments on LLaMA 3.1-8B, and additionally performed extra experiments on Mistral 7B-v0.3.
  • LLaMA 3.1-8B (Dubey et al., 2024): LLaMA 3.1-8B is a large language model released by Meta, featuring 8 billion (8B) parameters. It is part of the LLaMA (Large Language Model Meta AI) series, focusing on efficient reasoning and text generation capabilities. LLaMA 3.1-8B excels in code generation, language understanding, and conversational tasks, optimizing inference speed and training efficiency, making it suitable for research, commercial applications, and AI studies.
  • Mistral 7B-v0.3 (Jiang et al., 2023): Mistral 7B-v0.3 is an open-source language model developed by Mistral AI, featuring 7 billion parameters. It is optimized based on the Transformer architecture, emphasizing efficiency and multitasking capabilities. Compared to earlier versions, this model shows improvements in coding, mathematics, and reasoning tasks, making it suitable for chatbots, programming assistance, and natural language processing tasks. Mistral 7B-v0.3 incorporates feedback from the open-source community to enhance inference efficiency, delivering high performance with reduced computational resources.
C.3 Benchmarks
We assess model performance on general knowledge, mathematical problem-solving, code generation and commonsense reasoning benchmarks.
  • IFEval (Instruction Following Evaluation) (Zhou et al., 2023c): a benchmark dataset designed to assess the instruction-following capabilities of large models. It encompasses various tasks, including general knowledge question answering, commonsense reasoning, and mathematical reasoning, aiming to measure the understanding and accuracy of language models when executing complex instructions.
  • MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021a): a large-scale, multi-task language understanding benchmark that covers 57 subjects, testing models on their knowledge and reasoning abilities across fields such as history, law, mathematics, and medicine. It serves as a significant indicator of general artificial intelligence knowledge levels.
  • GSM8K (Grade School Math 8K) (Cobbe et al., 2021): a dataset specifically created for solving mathematical problems, containing approximately 8,500 elementary school math questions that primarily focus on basic arithmetic, logical reasoning, and text comprehension skills. This dataset is used to evaluate models’ mathematical computation and reasoning abilities.
  • MATH (Hendrycks et al., 2021b): consists of math competition problems from middle school and college levels, covering areas such as algebra, geometry, number theory, and calculus. This dataset is more challenging than GSM8K and is primarily used to assess models’ performance on advanced mathematical reasoning tasks.
  • HumanEval (Chen et al., 2021): a dataset for evaluating code generation capabilities, featuring a series of Python programming problems, each with a clear function signature and test cases. This dataset is commonly used to measure AI performance in automated code generation and programming tasks.
  • MBPP (Mostly Basic Programming Problems) (Austin et al., 2021): a benchmark dataset for code generation, containing 1,000 basic programming questions that cover data structures, algorithms, and logical reasoning. It is suitable for assessing AI capabilities in fundamental programming tasks.
  • Hellaswag (Zellers et al., 2019): a benchmark dataset for commonsense reasoning, consisting of a series of incomplete sentences that require models to select the most reasonable ending. This dataset tests models’ contextual understanding and reasoning abilities by designing misleading options.
  • GPQA (Graduate-Level Google-Proof Q&A) (Rein et al., 2024): a challenging dataset designed to evaluate the capabilities of LLMs and scalable oversight mechanisms. Let me provide more details about it.
C.4 Baselines
We compare Middo with both existing data selection and data augmentation methods on the Alpaca dataset.
Data Selection Methods.
  • Alpaca-clean (Ruebsamen, 2023): a cleaned version of the Alpaca dataset that removes low-quality samples and duplicates, aiming to improve the overall quality of the dataset.
  • Superfiltering (Li et al., 2024b): using smaller, weaker language models (such as GPT-2) as data filters to compute IFD allows for the selection of high-quality instruction tuning data.
  • Long (Zhao et al., 2024a): directly select the 1,000 samples with the longest responses as training data.
  • AlpaGasus (Chen et al., 2024): utilize powerful LLMs (such as ChatGPT) to automatically assess the sample quality in the Alpaca dataset and filter out high-quality data to enhance model training effectiveness.
Data Augmentation Methods.
  • Alpaca-GPT4 (Peng et al., 2023): a data augmentation method that uses GPT-4 to generate additional training data for the Alpaca dataset.
  • I-SHEEP (Anonymous, 2025b): a data augmentation method that uses a self-supervised learning approach to generate additional training data for the Alpaca dataset.
  • WizardLM (Xu et al., 2024): 70K data generated based on Evol-Instruct, which aims to generate more complex instruction data through a recursive evolutionary approach in order to improve the model’s reasoning and instruction comprehension.
C.5 Hyperparameters
Hyperparameter Value
LLaMA-3.1-8B
Learning Rate $2\times 10^{-5}$
Number of Epochs 11
Number of Devices 88
Per-device Batch Size 44
Gradient Accumulation Steps 88
Learning Rate Scheduler cosine
Warmup Ratio $0.03$
Max Sequence Length $4096$
Mistral-7B-v0.3
Learning Rate $1\times 10^{-5}$
Number of Epochs 11
Number of Devices 88
Per-device Batch Size 44
Gradient Accumulation Steps 88
Learning Rate Scheduler cosine
Warmup Ratio $0.03$
Max Sequence Length $4096$
Table 8: Hyperparameters used for fine-tuning.
Fine-tune.
For LLaMA-3.1-8B, we follow the Alpaca GitHub repository [^1] https://github.com/tatsu-lab/stanford_alpaca, setting the batch size to 32, the learning rate to $2\times 10^{-5}$, and the warmup ratio to 0.03. For Mistral-7B-v0.3, we adjust the learning rate to $1\times 10^{-5}$, as per official recommendations [^2] https://docs.mistral.ai/capabilities/finetuning. All the hyperparameters are detailed in Table 8.
Data Synthetic.
We use the OpenAI API to generate data by GPT-4o-mini, setting both temperature and top_p to 1.0 to guarantee diversity.
Hyperparameter Value
Pass@n $n=1$
Presence Penalty $0.0$
Frequency Penalty $0.0$
Repetition Penalty $1.0$
Temperature $0.0$
Top_p $1.0$
Top_k 1-1
Min_p $0.0$
Max Tokens $4096$
Min Tokens 0
Table 9: Hyperparameters used for evaluation.
Evaluation.
All benchmarks are conducted in zero-shot and we conducted the tests using the default configuration of OpenCompass. All the hyperparameters are detailed in Table 9.
All experiments are conducted on 8 $\times$ NVIDIA Tesla A100 GPUs about 50 GPU hours.
Appendix D Self-alignment Scores
(a) Alpaca dataset.
(b) Alpaca-4o-mini dataset.
(c) WizardLM dataset.
Figure 8: Self-alignment score evolution across iterations. The x-axis represents the self-alignment scores, while the y-axis shows the density of data points.
We provide detailed self-alignment score evolution across iterations on the Alpaca, Alpaca-4o-mini, and WizardLM datasets in Figure 8. These figures illustrate the dynamic evolution of self-alignment scores across iterations, highlighting the continuous improvement in dataset quality and alignment with model capabilities.
Appendix E Case Study
Complexity Optimization Prompt Template You are a powerful LLM with the task to simplify the given #Prompt# for weaker open source LLMs (e.g. LLaMA, Qwen). Since the original #Prompt# is hard for them to handle, you need to rewrite it into a simpler version that these weaker LLMs can handle or learn from more easily. Follow the steps below carefully. Step 1: Read the #Prompt# in detail. Suggest possible methods to make this prompt easier for weaker LLMs to handle or learn from, then list these methods as your #Methods List#. Step 2: Create a comprehensive plan to simplify the #Prompt# using several methods from your #Methods List#. Present your plan in a clear, step-by-step format. Step 3: Execute your plan to rewrite the #Prompt# into a simpler, more learnable version. You can change scenarios, contexts, or settings as needed. Your goal is to ensure that weaker LLMs learn from this prompt, rather than just memorizing an answer. Step 4: Finally, review your rewritten version for any problems. Present only the #Final Rewritten Prompt#, without any additional explanation. Please respond strictly in the following format: Step 1 #Methods List#: Step 2 #Plan#: Step 3 #Rewritten Prompt#: Step 4 #Final Rewritten Prompt#:
Overly Hard Data Example Find the mode of the following set of numbers: 23, 16, 22, 19, 24, 21
Optimize Steps Step 1 #Methods List#:
1. Simplify the language used in the prompt.
2. Break down the instructions into simpler steps. 3. Provide examples to clarify the concept of mode. 4. Use smaller numbers or more relatable contexts. 5. Ask the question in a more direct manner. Step 2 #Plan#:
1. Change complex terminology like “mode” to a more descriptive explanation.
2. Use a familiar context, like favorite fruits or animals. 3. Clearly outline the steps needed to find the most common number, potentially using a smaller set. 4. Offer a simple example to illustrate the mode calculation. 5. Keep the structure straightforward and direct. Step 3 #Rewritten Prompt#:
Find the most frequent number in this list: 3, 7, 2, 3, 5, 7. What number appears the most?’
Step 4 #Final Rewritten Prompt#:
Find the number that appears the most in this list: 3, 7, 2, 3, 5, 7. What is that number?
Optimized Easier Data Find the number that appears the most in this list: 3, 7, 2, 3, 5, 7. What is that number?
Figure 9: Complexity optimization example.
Diversity Extension Prompt Template You are a powerful LLM with the task to create brand new prompts for weaker open source LLMs (e.g. LLaMA, Qwen). You need to create a brand new complete prompt for them to learn in order to improve their knowledge and skills. Follow the steps below carefully. Use #Hint Prompt 1# and #Hint Prompt 2# as guiding examples. Then read the #Core Prompt# in detail. Be inspired to suggest additional new prompts, and ultimately create only one completely original and diverse #New Prompt#. Please respond strictly in the following format: #New Prompt#:"
Sparse Data And Neighbors #Hint Prompt 1#:
How long did Shakespeare live?
#Hint Prompt 2#:
How did the Industrial Revolution change society?
#Core Prompt#:
How did Julius Caesar die?
Extensioned Data What were the key factors that led to the fall of the Roman Empire?
Figure 10: Diversity extension example.
Quality Optimization Prompt Template You are a powerful LLM with the task to rewrite the given #Prompt# for weaker open source LLMs (e.g., LLaMA, Qwen). Since the original #Prompt# is of poor quality for them to learn, you need to rewrite it into a higher quality version that these weaker models can better learn from. Follow the steps below carefully. Step 1: Read the #Prompt# in detail. Identify reasons for the poor quality of the #Prompt#. Suggest possible methods to improve the quality of the #Prompt#, then list these methods as your #Methods List#. Step 2: Create a comprehensive plan to rewrite the #Prompt# using several methods from your #Methods List#. Present your plan in a clear, step-by-step format. Step 3: Execute your plan to rewrite the #Prompt# into a higher quality version. Step 4: Finally, review your rewritten version for any problems. Present only the #Final Rewritten Prompt#, without any additional explanation. Please respond strictly in the following format: Step 1 #Methods List#: Step 2 #Plan#: Step 3 #Rewritten Prompt#: Step 4 #Final Rewritten Prompt#:
Low Quality Data Example Given the following input, generate a job description for a product manager. Noinput
Optimize Steps Step 1 #Methods List#:
1. Clarify the input requirements to make them more explicit.
2. Use structured language with clear expectations for the output. 3. Specify the context in which the job description will be used. 4. Provide examples of desirable skills and experiences for the product manager role. 5. Ensure the language is professional and appropriate for a job description. Step 2 #Plan#:
1. Add a clear request for a job description that outlines key responsibilities and qualifications for a product manager.
2. Specify the context, such as whether it’s for a tech company, a startup, or a different industry. 3. Suggest including attributes and skills relevant to the product manager role, such as project management, communication, and analytics. 4. Format the request to clearly indicate the sections that should be included in the job description (e.g., summary, responsibilities, qualifications). Step 3 #Rewritten Prompt#:
<note:same as next box>
Step 4 #Final Rewritten Prompt#:
<note:same as next box>
Optimized High Quality Data Please create a detailed job description for a Product Manager position in a tech company. The description should include the following sections: - Job Summary: A brief overview of the role and its importance within the company. - Key Responsibilities: A list of primary duties that the Product Manager will be responsible for, such as managing product development, collaborating with cross-functional teams, and conducting market research. - Required Qualifications: Skills and experiences necessary for the role, including knowledge of product management methodologies, strong analytical abilities, and effective communication skills. - Preferred Attributes: Any additional qualities that would make a candidate stand out, such as a passion for technology or experience in agile environments.
Figure 11: Quality optimization example.
Self-Alignment Instruction Score Prompt Template We would like to request your feedback on the {dimention} of the prompt displayed below. Prompt: Give three tips for staying healthy. Please rate according to the {dimention} of the prompt to evaluate {explanations}. Each prompt is scored on a scale of 0 to 10, with higher scores indicating higher {dimention}. Try to avoid scoring a full 10. Give your rating number first, then give a explanation of your rating.
Self-Alignment Instruction Score Dimentions and Explanations "factuality": "whether the information provided in the prompt is accurate and based on reliable facts and data", "clarity": "whether the prompt is clear and understandable, and whether it uses concise language and structure", "completeness": "whether the prompt provides sufficient information and details".
Instruction Clarity Score Example 7. The prompt is clear and understandable, but it could be more concise.
Figure 12: Self-Alignment instruction score example.
Self-Alignment Response Score Prompt Template We would like to request your feedback on the {dimention} of the prompt displayed below. Prompt: What are the three primary colors? Response: The three primary colors are red, blue, and yellow. Please rate according to the {dimention} of the response to evaluate {explain}. Each response is scored on a scale of 0 to 10, with higher scores indicating higher {dimention}. Try to avoid scoring a full 10. Give your rating number first, then give a explanation of your rating.
Self-Alignment Response Score Dimentions and Explanations "factuality": "whether the information provided in the response is accurate and based on reliable facts and data", "clarity": "whether the response is clear and understandable, and whether it uses concise language and structure", "completeness": "whether the response provides sufficient information and details".
Response Clarity Score Example 8.5. The response is clear and understandable, but it could be more concise.
Figure 13: Self-Alignment response score example.

### v2

@@ 5 Analysis
5.1 Ablation Studies
To assess the effectiveness of Middo and the contribution of each optimization pipeline, we conduct ablation experiments with the LLaMA-3.1-8B model on the Alpaca dataset. Specifically, we analyze the following ablations: (a) w/o loss: removes Loss Patterns. (b) w/o neighbor: excludes Embedding Cluster Dynamics. (c) w/o score: removes Self-alignment Scores.
Iter. Ablations IFEval MATH HumanEval Hellaswag Avg.
@@ 5.2 Effect of Selected Data Scale
We investigate the impact of the different scales of the selected and optimized data in this section by varying the thresholds for data selection. Results are illustrated in Figure 3. We observe that increasing the size of the refined data initially leads to an upward trend in performance; however, once the refined data exceeds a certain threshold, performance begins to decline. To maintain the potential for further iterative improvement, we set the refined data size at a moderate level that optimally balances the cost and benefit of the optimization process. In the first iteration, each component selects approximately $5\%$ of the data for refinement. By controlling the parameter $m$, the amount of data refined can adaptively change as the model’s capability increases. Detailed data sizes selected in each iteration are provided in Appendix 11.
@@ 5.3 Data Analysis
Dynamic Iterative Improvement.
For a deeper understanding of how Middo transforms the dataset, we provide an analysis of its impact on data complexity, diversity, and quality.
@@
To quantify how Middo modulates dataset complexity, we analyze the loss distribution evolution through optimization cycles. As shown in Figure 4, the original dataset exhibits a long-tailed distribution with extreme loss values up to $12.99$. After applying Middo, the maximum loss decreases by $71.05\%$ to $3.76$, indicating successful mitigation of overly complex samples and the distribution mode shifts leftward, suggesting better alignment between data complexity and model capability. This transformation demonstrates our framework’s ability to adaptively prune pathological samples while preserving pedagogically valuable challenges.
@@
To analyze the diversity of the dataset after applying Middo, we visualize the data distribution using t-SNE [50]. Figure 5 reveals how the augmented data points are distributed relative to the original data. Notably, most of the augmented samples are located at the peripheries of the clusters, effectively filling in the sparsely populated regions. This distribution indicates that Middo is not merely adding redundant data but is instead enhancing the overall coverage of the latent space. By strategically augmenting the dataset at the cluster edges, Middo improves the diversity and ensures a more uniform distribution of data points, ultimately contributing to better model generalization.
@@
Figure 4: Loss distribution comparison before and after applying Middo. The density curve reflects the relative frequency of data points within specific loss intervals. The inset subfigure highlights the maximum loss reduction from $12.99$ to $3.76$.
Figure 5: t-SNE visualization of the Alpaca dataset before and after applying Middo. The original dataset is shown in light blue, while the augmented data is in dark blue. The dark blue points tend to occupy the sparsely populated regions of the light blue point distribution.
@@ 6 Conclusion
In this paper, we present Middo, a model-informed dynamic data optimization framework that transforms LLM fine-tuning via closed-loop learning. Unlike traditional static methods, Middo establishes a self-evolving system that continuously adapts to the model’s evolving capabilities. It employs three core mechanisms: complexity optimization refines overly complex samples using loss patterns, ensuring the training data remains appropriately challenging; diversity optimization enhances dataset diversity by analyzing embedding cluster dynamics; and quality optimization leverages self-alignment scores to evaluate and improve the quality of training samples. Experiments on multiple benchmarks demonstrate that Middo consistently boosts LLMs’ performance, achieving an average accuracy improvement of $7.15\%$ while maintaining the original data scale on LLaMA-3.1-8B. Ablation studies confirm the effectiveness of each component, underscoring the importance of balancing complexity, diversity, and quality. Middo ’s adaptability and model-awareness make it a powerful tool for sustainable LLM training. Moreover, our approach paves the way for future research in adaptive training that continuously optimizes learning efficiency.
@@
Acknowledgments
This work is supported by Shanghai Artificial Intelligence Laboratory. Zinan Tang is an intern at Shanghai Artificial Intelligence Laboratory.